OpenStudy (anonymous):

Let f(x)=4x+1 and g[f(x)]=(102-x)(x+998). Compute g(9) and g(99).

OpenStudy (anonymous):

At g(9), f(x)=9 which means \(4x+1=9 \implies x=2\). So \(g(9)=(102-2)(2+998)=..\)